OpenStudy (anonymous):

Use the Quadratic Formula to solve the equation. -2x ^2 - 5x + 5 = 0

OpenStudy (anonymous):

do you know formula or not

OpenStudy (anonymous):

@shopat just use this https://www.mathway.com/

OpenStudy (the_fizicx99):

A = -2 B = -5 C = 5 \(\ \sf x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \) \(\ \sf \dfrac{ -(-5) \pm \sqrt{-5^2 -4(-2)(5)}}{2(-2)} \) \(\ \sf \dfrac{5 \pm \sqrt{25 - 40}}{-4} \) Continue to solve.

OpenStudy (the_fizicx99):

Typo, \(\ \ \sf \dfrac{5 \pm \sqrt{25 + 40}}{-4} \)

OpenStudy (anonymous):

@Maria195 thank you!! @tHe_FiZiCx99 - I got it thanks!!

OpenStudy (the_fizicx99):

You're welcome!

OpenStudy (anonymous):

You're welcome :D